# The Mittag-Leffler theorem for proper minimal surfaces and directed meromorphic curves

### v1

@@ 5.  The Mittag-Leffler theorem for meromorphic $A$-immersions
In this section we establish the Mittag-Leffler theorem with approximation and interpolation for directed meromorphic immersions by proving the following more precise version of the first part of Theorem 1.6, including also condition (iv). We shall complete the proof of Theorem 1.6 by checking assertion  (v) and (vi), concerning global geometry, in Section 7 (see Theorem 7.1).
@@
    $G-F$ is continuous on $E$, that is, it extends continuously to $U$.
@@
    $20<\varepsilon_{j}<\varepsilon_{j-1}/$ 2, and if $G\colon M\setminus E\to\mathbb{C}^{n}$ is a holomorphic map with $||G-F^{j}||_{M_{j}\setminus(E\cap M_{j})}<2\varepsilon_{j}$, then $G$ is an immersion on $M_{j-1}\setminus(E\cap M_{j-1})$.
The base of induction is provided by the already chosen $\varepsilon_{0}$ and $F^{0}$. (Note that $M_{0}\cap U=\varnothing$ and $F^{0}\colon M_{0}\to\mathbb{C}^{n}$ is a full holomorphic $A$-embedding.) Properties (10) and (50) are void, whereas (20)–(40) hold for free as $U\cap M_{0}=\varnothing$. For the inductive step, assume that for some fixed $j\in\mathbb{N}$ we have already found numbers $\varepsilon_{i}>0$ and full meromorphic $A$-immersions $F^{i}$ satisfying (2i)–(4i) for all $i\in\{0,1,\dots,j-1\}$. Let us construct $\varepsilon_{j}$ and $F^{j}$. Firstly, we shall connect $M_{j-1}$ and $W_{j}\subset\mathring{M}_{j}\setminus M_{j-1}$ to obtain a compact connected admissible set $L_{j}$ which is Runge in $M$. For that, pick points $p_{j}\in bM_{j-1}$ and $q_{j}\in bW_{j}$ such that $q_{j}$ does not belong to any of the finitely many Jordan arcs in the description of $W_{j}$ as admissible set (see Definition 2.3). Take a smooth oriented arc $\Gamma_{j}\subset\mathring{M}_{j}$ with the endpoints $p_{j}$ and $q_{j}$ such that it intersects $M_{j-1}\cup W_{j}$ only at its endpoints $p_{j},q_{j}$ and the intersection is transverse. Apply [6, Lemma 3.5.4] and extend $f^{j-1}\colon M_{j-1}\setminus(E\cap M_{j-1})\to A_{*}$ smoothly to $\Gamma_{j}$ such that it still maps to $A_{*}$ and satisfies
@@
$G:=\lim_{j\to\infty}F^{j}\colon M\setminus E\to\mathbb{C}^{n}$
@@
holds for each $j\geq 1$, and, noting that $U\cap M_{0}=\varnothing$, we conclude that $||G-F||_{U}<\varepsilon$, as desired. The fullness of $G$ is achieved for $\varepsilon_{0}>0$ small enough. Finally, in view of (4j) and arguing as in the proof of Theorem 4.1, if $F|_{\Lambda}$ is injective, then we can grant that $G$ is an embedding by choosing each number $\varepsilon_{j}>0$ sufficiently small. ∎
@@ 6.  The Mittag-Leffler theorem for meromorphic $A$-immersions
    $||\widetilde{F}^{\prime\prime}-F^{\prime\prime}||_{U}<\varepsilon$.
@@
Assume that $M$, $S$, $E$, $A$ and $r$ are as in Proposition 3.2, $\Lambda\subset\mathring{S}\setminus E$ is a finite set, and $A^{\prime}$, $\pi$, $k$, $m$ and $n$ are as is Theorem 6.1. Let $F=(F^{\prime},F^{\prime\prime})\colon S\setminus E\to\mathbb{C}^{n}$ be a map of class $\mathscr{I}_{A}(S|E)$ and assume that $F^{\prime}$ extends to a nonconstant meromorphic map $F^{\prime}\colon M\to\mathbb{C}^{k}$ that is holomorphic on $M\setminus E$ and such that the zero set $(f_{1},\dots,f_{k})^{-1}(0,\dots,0)$ intersects $S$ only at points in $\mathring{S}=\mathring{K}$, where $F^{\prime}$, $F^{\prime\prime}$ and $f_{i}$ are defined as in Theorem 6.1. Given $\varepsilon>0$, there exists a meromorphic $A$-immersion $\widetilde{F}=(F^{\prime},\widetilde{F}^{\prime\prime})\in\mathscr{I}_{A}(M|E)$ satisfying the following.
@@
    $||\widetilde{F}^{\prime\prime}-F^{\prime\prime}||_{S}<\varepsilon$.
@@
We combine ideas from the proof of Theorem 3.1 with those of the proofs of [4, Theorem 7.7] and [6, Theorem 3.7.1]. Firstly, observe that by assumptions on $A^{\prime}=A\cap\{z_{1}=1,\dots,z_{k}=1\}$, the projection $\pi=(\pi_{1},\dots,\pi_{k})\colon A\to\mathbb{C}^{k}$ is a trivial fibre bundle with Oka fibre $A^{\prime}$ except over $(0,\dots,0)\in\mathbb{C}^{k}$. Given a nowhere vanishing holomorphic $1$-form $\theta$ on $M$, we write $dF=f\theta$ and $f=(f^{\prime},f^{\prime\prime})$ for $f^{\prime}=(f_{1},\dots,f_{k})$ and $f^{\prime\prime}=(f_{k+1},\dots,f_{n})$. By assumption, $f^{\prime}$ is meromorphic and nonconstant on $M$, hence its zero set $Q:=(f^{\prime})^{-1}(0)=\{q_{1},q_{2},\dots\}$ is discrete in $M$. Take a holomorphic function $g\in\mathscr{O}(M)$ with
@@
$\tilde{f}=(h^{\prime},\tilde{h}^{\prime\prime})/g:M\setminus E\to A_{*},$
the meromorphic 1-form $\tilde{f}\theta$ on $M$ integrates to a meromorphic $A$-immersion $\widetilde{F}\colon M\setminus E\to\mathbb{C}^{n}$ satisfying the proposition. In order to achieve that, we wish to apply [6, Theorem 1.13.4] (cf. [16, Theorem 6.14.6]); however, we need a globally defined continuous section of $\Sigma$ over $M$ which is holomorphic on a small neighbourhood of any point in $M$ over which $(h^{\prime})^{*}\pi$ is ramified, i.e., on $Q\cup E$, and on a small neighbourhood of any point of interpolation, i.e., on $\Lambda\cup E$. Note that if $p\in(Q\cup E\cup\Lambda)\cap S$, then $p\in\mathring{S}=\mathring{K}$ and $h^{\prime\prime}$ is holomorphic around $p$ by assumption. The other possibility is that $p\in Q\cap(M\setminus S)$, because $E\cup\Lambda\subset\mathring{S}$. So, on a disc neighbourhood of any such point $p$ we choose a holomorphic section $h^{\prime\prime}$ such that $h^{\prime\prime}(p)\neq 0$ and $(h^{\prime},h^{\prime\prime})$ maps to $A_{*}$. We add all these disc neighbourhoods to the domain of holomorphicity of $h^{\prime\prime}$.
We aim at approximating $h^{\prime\prime}$ on $S$ and interpolating on $\widetilde{\Lambda}:=Q\cup E\cup\Lambda$ at the same time. Observe that $(M,S\cup\widetilde{\Lambda})$ is homotopy equivalent to a relative CW-complex of dimension $1$, therefore, a skeleton of $M$ is obtained from $S\cup\widetilde{\Lambda}$ by attaching points and edges. Moreover, regular fibres $A^{\prime}_{c_{1},\dots,c_{k}}:=A\cap\{z_{1}=c_{1},\dots,z_{k}=c_{k}\}$ for $(c_{1},\dots,c_{k})\neq(0,\dots,0)$ are connected, which implies that any holomorphic section of $\Sigma$ over $S\cup\widetilde{\Lambda}$ can be extended to a continuous section over $M$. Hence, the assumptions in [6, Theorem 1.13.4] (the basic Oka principle for sections of ramified maps) are satisfied. We shall adapt the proof of Lemma 3.3 to the current aim.
@@
(6.2) $\mathscr{P}=\left(\mathscr{P}_{1},\dots,\mathscr{P}_{l}\right)\colon\Big{\{}\hat{h}^{\prime\prime}\in\mathscr{C}(C,\mathbb{C}^{m})\colon\frac{\hat{h}^{\prime\prime}}{g}-f^{\prime\prime}\in\mathscr{C}^{0}(C,\mathbb{C}^{m})\Big{\}}\to\left(\mathbb{C}^{m}\right)^{l},$
@@
(6.4) $\Phi_{h}:=\left(\Phi_{h^{\prime}},\Phi_{h^{\prime\prime}}\right)\colon U\times S\to\mathbb{C}^{n}$
@@
Step 3: The noncritical case. Assume that $S$ is a connected admissible set in an open Riemann surface $M$ such that the inclusion $S\hookrightarrow M$ induces an isomorphism $H_{1}(S,\mathbb{Z})\to H_{1}(M,\mathbb{Z})$. Let $\widetilde{\Lambda}\subset S$ be a finite subset and $\mathscr{C}$ be as above. (Note that at the moment, $S$ and $\widetilde{\Lambda}$ are not the ones in the statement of the proposition.) Assume that $(h^{\prime})^{*}\pi$ is as above, with its ramification values in $\mathring{S}=\mathring{K}$, and $h=(h^{\prime},h^{\prime\prime})$ and $\Phi_{h}$ are as in the previous steps for this pair $(M,S)$. By [6, Theorem 1.13.1, Theorem 1.13.4], there exists a holomorphic section $\hat{h}^{\prime\prime}$ of $\Sigma$ on $M$ which approximates $h^{\prime\prime}$ uniformly on $S$, and is such that $\hat{h}^{\prime\prime}/g-f^{\prime\prime}$ is continuous on $S$ and $\hat{h}^{\prime\prime}-h^{\prime\prime}$ vanishes to order $r^{\prime}$ on $\widetilde{\Lambda}$. Replace $h=(h^{\prime},h^{\prime\prime})$ by $\hat{h}=(h^{\prime},\hat{h}^{\prime\prime})\colon M\to A_{*}$ as the core of the spray $\Phi_{h}$ in (6.4), then interpolate and approximate uniformly on $S$ all functions used in the construction of $\Phi_{h}$ by global holomorphic functions on $M$ (see the proof of Claim 3.5). The new spray $\Phi_{\hat{h}}\colon\widehat{U}\times M\to A_{*}$ (where $0\in\widehat{U}\subset U\subset\mathbb{C}^{ml}$), given by $\Phi_{\hat{h}}=(\Phi_{h^{\prime}},\Phi_{\hat{h}^{\prime\prime}})$, has the same properties as $\Phi_{h}$ provided that all approximations are close enough. Consequently, the implicit function theorem guarantees existence of a point $t_{0}\in\widehat{U}$ near the origin such that the map $\tilde{h}^{\prime\prime}:=\Phi_{\hat{h}^{\prime\prime}}(t_{0},\cdot)$ is a section of $M\to\Sigma$ satisfying $\mathscr{P}(\tilde{h}^{\prime\prime})=0,\ \tilde{h}^{\prime\prime}\approx h^{\prime\prime}$ on $S$, $\tilde{h}^{\prime\prime}/g-f^{\prime\prime}$ is continuous on $S$ and $\tilde{h}^{\prime\prime}-h^{\prime\prime}$ vanishes to order $r^{\prime}$ on $\widetilde{\Lambda}$.
@@
Step 4: The induction and the critical case. We proceed by induction as in the proof of Proposition 3.2. We shall briefly explain the procedure at the $j$-th step ($j\geq 1$) where we have a suitable meromorphic section $(h^{\prime\prime})^{j-1}$ of $\Sigma$ defined on the compact set $M_{j-1}$ and wish to approximate it with interpolation by a section $(h^{\prime\prime})^{j}$ defined on $M_{j}$ (here, the sets $M_{j}$ form an exhaustion of $M$ as in the proof of Proposition 3.2). We distinguish two cases. In the first one, the annulus $A_{j}$ (defined as in the proof of Proposition 3.2) does not contain any critical points of the Morse exhaustion function $\tau$ on $M$; this case is solved by the noncritical case. On the contrary, when $p\in A_{j}$ is the unique critical point of $\tau$ on $A_{j}$, as described in the proof of the cited proposition, we add certain arcs (denoted by $\Gamma_{j}$) if necessary. We keep the first $k$ component functions fixed (i.e., $h^{\prime}=(h^{\prime})^{j}$), and apply the argument to $(h^{\prime\prime})^{j-1}$, such that the extension of $(h^{\prime\prime})^{j-1}$ to $\Gamma_{j}$ is a holomorphic section of $\Sigma$ and $\int_{\Gamma_{j}}\frac{(h^{\prime\prime})^{j-1}}{g}\theta$ equals a prescribed vector in $\mathbb{C}^{m}$. (Note that the use of an analogue of Gromov integration lemma is justified by [6, Remark 3.5.5]).
@@
Step 5. Completion of the proof. Performing the induction process, we then obtain in the limit a holomorphic map $\tilde{h}^{\prime\prime}=\lim_{j\to\infty}(h^{\prime\prime})^{j}:M\to\mathbb{C}^{m}$, such that $(h^{\prime},\tilde{h}^{\prime\prime}):M\to\mathbb{C}^{n}$ takes its values in $A_{*}$, $\tilde{h}^{\prime\prime}/g$ is holomorphic on $M\setminus E$, the difference $\tilde{h}^{\prime\prime}-h^{\prime\prime}$ is uniformly close to 00 on $S$ and it vanishes to order $r^{\prime}$ at every point in $E\cup\Lambda$. Thus, $\tilde{h}^{\prime\prime}/g-f^{\prime\prime}$ is continuous on $S$ and vanishes to order $r$ everywhere on $E\cup\Lambda$. Furthermore, by the control of periods, we have that $(\tilde{h}^{\prime\prime}/g)\theta$ is exact on $M\setminus E$ and the holomorphic map $\widetilde{F}^{\prime\prime}:M\setminus E\to\mathbb{C}^{m}$, defined by
@@
for any given base point $p_{0}\in S\setminus E$, satisfies that $\widetilde{F}^{\prime\prime}-F^{\prime\prime}$ is continuous and $\varepsilon$-close to 00 on $S$, and it vanishes to order $r$ at every point in $E\cup\Lambda$. It is now clear that the well-defined meromorphic $A$-immersion $\widetilde{F}:=(F^{\prime},\widetilde{F}^{\prime\prime}):M\setminus E\to\mathbb{C}^{n}$ of class $\mathscr{I}_{A}(M|E)$ satisfies the conclusion of the proposition. ∎
@@
Let $M$, $A$ and $n$ be as in Theorem 7.1 (i), assume that $K$ and $L$ are smoothly bounded Runge compact domains in $M$ such that $K\subset\mathring{L}$, and $E,\Lambda\subset\mathring{K}$ are finite disjoint sets. Let $F\colon K\setminus E\to\mathbb{C}^{n}$ be a meromorphic $A$-immersion of class $\mathscr{I}_{A}(K|E)$. Then for any numbers $r\in\mathbb{N}$, $\varepsilon>0$ and $\delta>0$, there is a meromorphic $A$-immersion $\widetilde{F}=(\widetilde{F}_{1},\widetilde{F}_{2},\ldots,\widetilde{F}_{n})\in\mathscr{I}_{A}(L|E)$ such that the following hold.
@@
Let $M$, $A$ and $n$ be as in Theorem 7.1 (ii), and assume that $K$, $L$, $E$ and $\Lambda$ are as in Lemma 7.2. Let $F\colon K\setminus E\to\mathbb{C}^{n}$ be a meromorphic $A$-immersion of class $\mathscr{I}_{A}(K|E)$ mapping $bK$ into $\mathbb{C}^{n}\setminus\{0\}$. Let $r\in\mathbb{N}$ and $\varepsilon>0$, and pick a number
$0<\delta<\min\left\{||F(p)||_{\infty}\colon p\in bK\right\},$
@@
    $||\widetilde{F}(p)||_{\infty}>\delta$ for all $p\in L\setminus\mathring{K}$.
@@
    $||\widetilde{F}(p)||_{\infty}>1/\varepsilon$ for all $p\in bL$.
The proof is very similar to that of [6, Lemma 3.11.1] (see also [4, Lemma 8.2]) and we leave the details out. The main difference is that we use Propositions 3.2 and 6.2 in place of [6, Theorems 3.6.1 and 3.7.1]. Moreover, in the inductive process, we ask the map $F^{b}$, constructed at the $b$-th step of induction, to satisfy analogous properties to (C1b)–(C7b) except (C3b) in the cited proof, and also that the difference $F^{b}-F^{b-1}$ is continuous on $L$ and vanishes to order $r$ on $E\cup\Lambda$, which is possible by our mentioned results.

### v2

@@ 5. The Mittag-Leffler theorem for meromorphic $A$-immersions
In this section we establish the Mittag-Leffler theorem with approximation and interpolation for directed meromorphic immersions by proving the following more precise version of the first part of Theorem 1.6, including also condition (iv). We shall complete the proof of Theorem 1.6 by checking assertions  (v) and (vi), concerning global geometry, in Section 7 (see Theorem 7.1).
@@
    GFG-F is continuous on $E$, that is, it extends continuously to $U$.
@@
    $0<\varepsilon_{j}<\varepsilon_{j-1}/2$, and if $G\colon M\setminus E\to\mathbb{C}^{n}$ is a holomorphic map with $G-F^{j}||_{M_{j}\setminus(E\cap M_{j})}<2\varepsilon_{j}$, then $G$ is an immersion on $M_{j-1}\setminus(E\cap M_{j-1})$.
The base of induction is provided by the already chosen $\varepsilon_{0}$ and $F^{0}$. (Note that $M_{0}\cap U=\varnothing$ and $F^{0}\colon M_{0}\to\mathbb{C}^{n}$ is a full holomorphic $A$-embedding.) Properties (10) and (50) are void, whereas (20)–(40) hold for free as $U\cap M_{0}=\varnothing$. For the inductive step, assume that for some fixed $j\in\mathbb{N}$ we have already found numbers $\varepsilon_{i}>0$ and full meromorphic $A$-immersions $F^{i}$ satisfying (2i)–(4i) for all $i\in\{0,1,\dots,j-1\}$. Let us construct $\varepsilon_{j}$ and $F^{j}$. Firstly, we shall connect $M_{j-1}$ and $W_{j}\subset\mathring{M}_{j}\setminus M_{j-1}$ to obtain a compact connected admissible set $L_{j}$ which is Runge in $M$. For that, pick points $p_{j}\in bM_{j-1}$ and $q_{j}\in bW_{j}$ such that $q_{j}$ does not belong to any of the finitely many Jordan arcs in the description of $W_{j}$ as admissible set (see Definition 2.3). Take a smooth oriented arc $\Gamma_{j}\subset\mathring{M}_{j}$ with the endpoints $p_{j}$ and $q_{j}$ such that it intersects $M_{j-1}\cup W_{j}$ only at its endpoints $p_{j},q_{j}$ and the intersection is transverse. Apply [7, Lemma 3.5.4] and extend $f^{j-1}\colon M_{j-1}\setminus(E\cap M_{j-1})\to A_{*}$ smoothly to $\Gamma_{j}$ such that it still maps to $A_{*}$ and satisfies
@@
$G:=\lim_{j\to\infty}F^{j}\colon M\setminus E\longrightarrow\mathbb{C}^{n}$
@@
holds for each $j\geq 1$, and, noting that $U\cap M_{0}=\varnothing$, we conclude that $G-F||_{U}<\varepsilon$, as desired. The fullness of $G$ is achieved for $\varepsilon_{0}>0$ small enough. Finally, in view of (4j) and arguing as in the proof of Theorem 4.1, if $F|_{\Lambda}$ is injective, then we can grant that $G$ is an embedding by choosing each number $\varepsilon_{j}>0$ sufficiently small. ∎
@@ 6. The Mittag-Leffler theorem for meromorphic $A$-immersions
    $\widetilde{F}^{\prime\prime}-F^{\prime\prime}||_{U}<\varepsilon$.
@@
Assume that $M$, $S$, $E$, AA and $r$ are as in Proposition 3.2, $\Lambda\subset\mathring{S}\setminus E$ is a finite set, and $A^{\prime}$, $\pi$, $k$, $m$ and $n$ are as is Theorem 6.1. Let $F=(F^{\prime},F^{\prime\prime})\colon S\setminus E\to\mathbb{C}^{n}$ be a map of class $\mathscr{I}_{A}(S|E)$ and assume that $F^{\prime}$ extends to a nonconstant meromorphic map $F^{\prime}\colon M\to\mathbb{C}^{k}$ that is holomorphic on $M\setminus E$ and such that the zero set $(f_{1},\dots,f_{k})^{-1}(0,\dots,0)$ intersects $S$ only at points in $\mathring{S}=\mathring{K}$, where $^{\prime}, F′′F^{\prime\prime}$ and $f_{i}$ are defined as in Theorem 6.1. Given $\varepsilon>0$, there exists a meromorphic $A$-immersion $\widetilde{F}=(F^{\prime},\widetilde{F}^{\prime\prime})\in\mathscr{I}_{A}(M|E)$ satisfying the following.
@@
    $\widetilde{F}^{\prime\prime}-F^{\prime\prime}||_{S}<\varepsilon$.
@@
We combine ideas from the proof of Theorem 3.1 with those of the proofs of [4, Theorem 7.7] and [7, Theorem 3.7.1]. Firstly, observe that by assumptions on $A^{\prime}=A\cap\{z_{1}=1,\dots,z_{k}=1\}$, the projection $\pi=(\pi_{1},\dots,\pi_{k})\colon A\to\mathbb{C}^{k}$ is a trivial fibre bundle with Oka fibre $A^{\prime}$ except over $(0,\dots,0)\in\mathbb{C}^{k}$. Given a nowhere vanishing holomorphic $1$-form $\theta$ on $M$, we write $dF=f\theta$ and $f=(f^{\prime},f^{\prime\prime})$ for $f^{\prime}=(f_{1},\dots,f_{k})$ and $f^{\prime\prime}=(f_{k+1},\dots,f_{n})$. By assumption, $f^{\prime}$ is meromorphic and nonconstant on $M$, hence its zero set $Q:=(f^{\prime})^{-1}(0)=\{q_{1},q_{2},\dots\}$ is discrete in $M$. Take a holomorphic function $g\in\mathscr{O}(M)$ with
@@
$\tilde{f}=(h^{\prime},\tilde{h}^{\prime\prime})/g:M\setminus E\longrightarrow A_{*},$
the meromorphic 1-form $\tilde{f}\theta$ on $M$ integrates to a meromorphic $A$-immersion $\widetilde{F}\colon M\setminus E\to\mathbb{C}^{n}$ satisfying the proposition. In order to achieve that, we wish to apply [7, Theorem 1.13.4] (cf. [24, Theorem 6.14.6]); however, we need a globally defined continuous section of $\Sigma$ over $M$ which is holomorphic on a small neighbourhood of any point in $M$ over which $(h^{\prime})^{*}\pi$ is ramified, i.e., on $Q\cup E$, and on a small neighbourhood of any point of interpolation, i.e., on $\Lambda\cup E$. Note that if $p\in(Q\cup E\cup\Lambda)\cap S$, then $p\in\mathring{S}=\mathring{K}$ and $h^{\prime\prime}$ is holomorphic around $p$ by assumption. The other possibility is that $p\in Q\cap(M\setminus S)$, because $E\cup\Lambda\subset\mathring{S}$. So, on a disc neighbourhood of any such point $p$ we choose a holomorphic section $h^{\prime\prime}$ such that $h^{\prime\prime}(p)\neq 0$ and $(h^{\prime},h^{\prime\prime})$ maps to $A_{*}$. We add all these disc neighbourhoods to the domain of holomorphicity of $h^{\prime\prime}$.
[…]
@@
(6.2) $\mathscr{P}=\left(\mathscr{P}_{1},\dots,\mathscr{P}_{l}\right)\colon\Big\{\hat{h}^{\prime\prime}\in\mathscr{C}(C,\mathbb{C}^{m})\colon\frac{\hat{h}^{\prime\prime}}{g}-f^{\prime\prime}\in\mathscr{C}^{0}(C,\mathbb{C}^{m})\Big\}\longrightarrow\left(\mathbb{C}^{m}\right)^{l},$
@@
(6.4) $\Phi_{h}:=\left(\Phi_{h^{\prime}},\Phi_{h^{\prime\prime}}\right)\colon U\times S\longrightarrow\mathbb{C}^{n}$
@@
Step 3: The noncritical case. Assume that $S$ is a connected admissible set in an open Riemann surface $M$ such that the inclusion $S\hookrightarrow M$ induces an isomorphism $H_{1}(S,\mathbb{Z})\to H_{1}(M,\mathbb{Z})$. Let $\widetilde{\Lambda}\subset S$ be a finite subset and $\mathscr{C}$ be as above. (Note that at the moment, $S$ and $\widetilde{\Lambda}$ are not the ones in the statement of the proposition.) Assume that $(h^{\prime})^{*}\pi$ is as above, with its ramification values in $\mathring{S}=\mathring{K}$, and $h=(h^{\prime},h^{\prime\prime})$ and $\Phi_{h}$ are as in the previous steps for this pair $(M,S)$. By [7, Theorem 1.13.1, Theorem 1.13.4], there exists a holomorphic section $\hat{h}^{\prime\prime}$ of $\Sigma$ on $M$ which approximates $h^{\prime\prime}$ uniformly on $S$, and is such that $\hat{h}^{\prime\prime}/g-f^{\prime\prime}$ is continuous on $S$ and $\hat{h}^{\prime\prime}-h^{\prime\prime}$ vanishes to order $r^{\prime}$ on $\widetilde{\Lambda}$. Replace $h=(h^{\prime},h^{\prime\prime})$ by $\hat{h}=(h^{\prime},\hat{h}^{\prime\prime})\colon M\to A_{*}$ as the core of the spray $\Phi_{h}$ in (6.4), then interpolate and approximate uniformly on $S$ all functions used in the construction of $\Phi_{h}$ by global holomorphic functions on $M$ (see the proof of Claim 3.5). The new spray $\Phi_{\hat{h}}\colon\widehat{U}\times M\to A_{*}$ (where $0\in\widehat{U}\subset U\subset\mathbb{C}^{ml}$), given by $\Phi_{\hat{h}}=(\Phi_{h^{\prime}},\Phi_{\hat{h}^{\prime\prime}})$, has the same properties as $\Phi_{h}$ provided that all approximations are close enough. Consequently, the implicit function theorem guarantees existence of a point $t_{0}\in\widehat{U}$ near the origin such that the map $\tilde{h}^{\prime\prime}:=\Phi_{\hat{h}^{\prime\prime}}(t_{0},\cdot)$ is a section of $M\to\Sigma$ satisfying $\mathscr{P}(\tilde{h}^{\prime\prime})=0,\ \tilde{h}^{\prime\prime}\approx h^{\prime\prime}$ on $S$, $\tilde{h}^{\prime\prime}/g-f^{\prime\prime}$ is continuous on $S$ and $\tilde{h}^{\prime\prime}-h^{\prime\prime}$ vanishes to order $r^{\prime}$ on $\widetilde{\Lambda}$.
@@
Step 4: The induction and the critical case. We proceed by induction as in the proof of Proposition 3.2. We shall briefly explain the procedure at the $j$-th step ($j\geq 1$) where we have a suitable meromorphic section $(h^{\prime\prime})^{j-1}$ of $\Sigma$ defined on the compact set $M_{j-1}$ and wish to approximate it with interpolation by a section $(h^{\prime\prime})^{j}$ defined on $M_{j}$ (here, the sets $M_{j}$ form an exhaustion of $M$ as in the proof of Proposition 3.2). We distinguish two cases. In the first one, the annulus $A_{j}$ (defined as in the proof of Proposition 3.2) does not contain any critical points of the Morse exhaustion function $\tau$ on $M$; this case is solved by the noncritical case. On the contrary, when $p\in A_{j}$ is the unique critical point of $\tau$ on $A_{j}$, as described in the proof of the cited proposition, we add certain arcs (denoted by $\Gamma_{j}$) if necessary. We keep the first $k$ component functions fixed (i.e., $h^{\prime}=(h^{\prime})^{j}$), and apply the argument to $(h^{\prime\prime})^{j-1}$, such that the extension of $(h^{\prime\prime})^{j-1}$ to $\Gamma_{j}$ is a holomorphic section of $\Sigma$ and $\int_{\Gamma_{j}}\frac{(h^{\prime\prime})^{j-1}}{g}\theta$ equals a prescribed vector in $\mathbb{C}^{m}$. (Note that the use of an analogue of Gromov integration lemma is justified by [7, Remark 3.5.5]).
@@
Step 5. Completion of the proof. Performing the induction process, we then obtain in the limit a holomorphic map $\tilde{h}^{\prime\prime}=\lim_{j\to\infty}(h^{\prime\prime})^{j}:M\to\mathbb{C}^{m}$, such that $(h^{\prime},\tilde{h}^{\prime\prime}):M\to\mathbb{C}^{n}$ takes its values in $A_{*}$, $\tilde{h}^{\prime\prime}/g$ is holomorphic on $M\setminus E$, the difference $\tilde{h}^{\prime\prime}-h^{\prime\prime}$ is uniformly close to 0 on $S$ and it vanishes to order $r^{\prime}$ at every point in $E\cup\Lambda$. Thus, $\tilde{h}^{\prime\prime}/g-f^{\prime\prime}$ is continuous on $S$ and vanishes to order $r$ everywhere on $E\cup\Lambda$. Furthermore, by the control of periods, we have that $(\tilde{h}^{\prime\prime}/g)\theta$ is exact on $M\setminus E$ and the holomorphic map $\widetilde{F}^{\prime\prime}:M\setminus E\to\mathbb{C}^{m}$, defined by
@@
for any given base point $p_{0}\in S\setminus E$, satisfies that $\widetilde{F}^{\prime\prime}-F^{\prime\prime}$ is continuous and $\varepsilon$-close to 0 on $S$, and it vanishes to order $r$ at every point in $E\cup\Lambda$. It is now clear that the well-defined meromorphic $A$-immersion $\widetilde{F}:=(F^{\prime},\widetilde{F}^{\prime\prime}):M\setminus E\to\mathbb{C}^{n}$ of class $\mathscr{I}_{A}(M|E)$ satisfies the conclusion of the proposition. ∎
@@
Let $M$, AA and $n$ be as in Theorem 7.1-(i), assume that $K$ and $L$ are smoothly bounded Runge compact domains in $M$ such that $K\subset\mathring{L}$, and $E,\Lambda\subset\mathring{K}$ are finite disjoint sets. Let $F\colon K\setminus E\to\mathbb{C}^{n}$ be a meromorphic $A$-immersion of class $\mathscr{I}_{A}(K|E)$. Then for any numbers $r\in\mathbb{N}$, $\varepsilon>0$ and $\delta>0$, there is a meromorphic $A$-immersion $\widetilde{F}=(\widetilde{F}_{1},\widetilde{F}_{2},\ldots,\widetilde{F}_{n})\in\mathscr{I}_{A}(L|E)$ such that the following hold.
@@
Let $M$, AA and $n$ be as in Theorem 7.1-(ii), and assume that $K$, $L$, $E$ and $\Lambda$ are as in Lemma 7.2. Let $F\colon K\setminus E\to\mathbb{C}^{n}$ be a meromorphic $A$-immersion of class $\mathscr{I}_{A}(K|E)$ mapping $bK$ into $\mathbb{C}^{n}\setminus\{0\}$. Let $r\in\mathbb{N}$ and $\varepsilon>0$, and pick a number
[…]
@@
    $\widetilde{F}(p)||_{\infty}>\delta$ for all $p\in L\setminus\mathring{K}$.
@@
    $\widetilde{F}(p)||_{\infty}>1/\varepsilon$ for all $p\in bL$.
The proof is very similar to that of [7, Lemma 3.11.1] (see also [4, Lemma 8.2]) and we leave the details out. The main difference is that we use Propositions 3.2 and 6.2 in place of [7, Theorems 3.6.1 and 3.7.1]. Moreover, in the inductive process, we ask the map $F^{b}$, constructed at the $b$-th step of induction, to satisfy analogous properties to (C1b)–(C7b) except (C3b) in the cited proof, and also that the difference $F^{b}-F^{b-1}$ is continuous on $L$ and vanishes to order $r$ on $E\cup\Lambda$, which is possible by our mentioned results.